# A Lensing-Band Approach to Spacetime Constraints

### 11 Dec 2023 v1

@@ IV.1.3 A seven-parameter family obtained from the leading-order terms in the KRZ expansion
As detailed above, we focus on the KRZ expansion (see Section IV.1.2) of circular spacetimes (see Section IV.1.1). Furthermore, we make sure to maintain oblate spheroidal coordinates at radial asymptotic infinity and match the leading order asymptotics to the mass $M$, the angular momentum $J$, and the PPN parameters $\beta$ and $\gamma$. The respective explicit derivation within the KRZ expansion is detailed in Appendix B. These choices result in a seven-parameter family of spacetimes for which the metric functions in Eq. 7 can be written in the following simple form
@@
We use the common shorthands $A=J/M$, $\Sigma=r^{2}+A^{2}\cos^{2}\theta$, and $\Delta=r^{2}-2Mr+A^{2}$ which are also often used to denote Kerr spacetime. In addition, we use the shorthand
@@
  • The parameter $a$ accounts for independent deviations of the horizon area [^11] In the context of the KRZ expansion, see Appendix B, $a$ is identified with the spin-parameter of the background Kerr spacetime. away from the horizon area of a Kerr black hole with equatorial horizon radius $r_{0}$ and asymptotic spin $A$.
@@ IV.2 From size constraints to shape constraints
Figure 6: An application of the lensing-band framework to the seven free parameters of the beyond-GR spacetime given in Eq. 8. Assuming the observation of a VLBI feature, as discussed in Section II.2, these projected constraints, to one standard deviation, result by demanding it to be completely within the first lensing band of the spacetime considered.
Given the spacetime in Eq. 8 and the lensed-emission region in Section II.2, we employ the lensing-band framework to obtain projected constraints on the underlying free parameters of the circular beyond-Kerr geometry, cf. Eq. 10.
The full results of the Bayesian parameter estimation, including deviations of all 7 free parameters $\Xi=(r_{0},M,A,\beta,\gamma,a,a_{01})$ of the underlying spacetime, is presented as corner plot in Fig. 6. As in the previous section, we choose flat priors on all the parameters with the following bounds: $0.01\leq r_{0}\leq 5$, $-5\leq A/M\leq 5$, $-1\leq\beta\leq 3$, $-1\leq\gamma\leq 5$, $-1\leq a\leq 1$, and $-5\leq a_{01}\leq 5$. As we will see, these flat priors are deformed into preferred regions, for most of these parameters, demonstrating the constraining power of a confident observation of a lensed emission region such as the one discussed in Section II.2.
We discuss several aspects of these results in more detail in the following section. To aid the discussion, we present explicit lensing bands of respective metric deformations and refer to the marginalized posteriors collectively presented in Fig. 6.
@@ IV.2.2 The effects of the inclination
As discussed in Section III, for the Kerr spacetime, the definition of equatorial lensing bands is ill-defined at edge-on inclination (i.e., at $i=90^{\circ}$), see, e.g., Fig. 7 in [44]. Hence, we are cautious to interpret results at close to edge-on inclination.
With the above caveat, and similar to the Kerr case in Section III, our results suggest a preference for low, i.e., close to face-on, inclination. Specifically, we find $i\doteqdot{32^{\circ}}^{+30^{\circ}}_{-22^{\circ}}\;$ which remains very similar to the posterior for the inclination obtained for the Kerr benchmark test in Section III. Once more, the lensing band becomes very thin at high, i.e., close to edge-on, inclination. Hence, it becomes increasingly delicate (or at some point impossible) to fit the observed lensed emission region into the theoretical lensing band. We re-iterate that such projected constraints formally rely on a confident observation of lensed emission everywhere within the associated image region.
@@ IV.2.3 Lifting (part of) the degeneracy among spherically symmetric PPN corrections $(\beta,\gamma)$
The PPN corrections ($\beta$ and $\gamma$) enter as spherically-symmetric deformations. Both PPN corrections remain consistent with the values preferred by GR ($\beta_{\text{GR}}=1$ and $\gamma_{\text{GR}}=1$). Moreover, they are correlated which suggests that the strongest projected constraint can actually be placed on the combination $(\beta-\gamma)$, see Fig. 6. This combination enters in the considered spacetime metric via $\Delta_{\beta\gamma}$ and, therefore, affects the $g_{tt}$ component of the metric, see Eq. 8. Several other works have quantified constraints directly in terms of constraints on the functional form of $g_{tt}$ [46, 37, 116, 60]. However, away from spherical symmetry or away from the critical curve, the lensing behaviour is not only controlled by the $g_{tt}$ component of the metric.
@@
In addition to this broadening/thinning effect, deformations of $\beta$ or $\gamma$, also lead to an increase/decrease of the overall size of the lensing band. Such a deformation in the overall size occurs generically for many spherically-symmetric deformation parameters and has previously been discussed in [37]. In fact, most constraints placed on deviations from GR to date reduce the geometrical information to the apparent size of the photon ring [47, 55] and are thus, as such, highly degenerate [37]. The degeneracy with the mass $M$, can, of course, be lifted by means of an external mass measurement [13]. However, without resolving geometric information beyond the overall apparent shadow diameter, degeneracies among various possible spherically-symmetric deformation parameters are impossible to avoid. Our results demonstrate that the first lensing band provides an independent way to break (at least some of these) these degeneracies. In particular, the lensing-band framework captures more information than just one overall diameter. Even in spherical symmetry, it captures continuous information encoded in the width of the lensing band. As a result, these degeneracies can (at least partially) be lifted. For instance, the independent lower bounds on $\beta$ and $\gamma$ indicates such a breaking of degeneracy.
@@ IV.2.4 Beyond PPN corrections: strong-field modifications
We have chosen to include deformations in $a_{01}$ in order to compare with results in the literature [47, 48, 55]. These studies assume black-hole uniqueness and therefore focus on $a_{01}$ and other higher-order deviation parameters. In an expansion around radial asymptotic infinity, $a_{01}$ contributes to the next-to-leading order spherically-symmetric deviations once the leading-order PPN parameters (i.e., $\beta$ and $\gamma$) have been fixed to their GR values. The latter are well-constrained at Solar System scales, while the former is not. In this sense, if one assumes black-hole uniqueness, then $a_{01}$ captures the leading deviations which have not already been constrained by other observations.
We find that the lensing-band framework places an independent projected constraint on $a_{01}$, both from above and from below, e.g, for the considered case $a_{01}\doteqdot-0.03^{+1.58}_{-1.22}$. Together with the theoretical bounds (see Appendix D), the lensing-band shape itself seems to be sufficient to break the degeneracy with all other deviation parameters investigated here. On the one hand, a theoretical constraint on the existence of a Killing horizon (cf. Appendix D) bounds the value of $a_{01}$ from below [56]. On the other hand, larger values of $a_{01}$ lead to a thinner lensing band, cf. Fig. 8, hence, eventually to an upper bound. Contrary to previous results [47, 48, 55], the obtained projected constraint does not rely on an external mass measurement. The projected constraint on $a_{01}$ has the same order of magnitude than the results inferred from the EHT observation of $*{}^{*}$ [33].
@@ IV.2.5 A shape constraint on the asymptotic angular momentum $J$
The lensing-band formalism also provides access to constraints which rely on the shape and not just on the overall size or width of the lensing band. In the following, we refer to these as “shape constraints”. Corresponding variations in the shape of the critical curve and in the second lensing band have been previously considered in Refs. [24, 41]. In particular, deformations of their approximately elliptical shape have lead to a proposed null test of GR [41, 45], see also Ref. [49] for beyond-GR deformations.
An explicit example of such a shape constraint is provided by deformations in the asymptotic angular momentum $J$, see Fig. 9. At large angular momentum, exceeding the viable values in Kerr spacetime, the lensing band shape is deformed so far that the lensed emission region modelled in Section II.2 can no longer be covered by the lensing band. Thus, the underlying geometry is excluded.
@@ V Discussion
We have introduced a framework for leveraging the observation of a lensed image of astrophysical emission surrounding a supermassive black hole to constrain the underlying spacetime geometry systematically. The approach is based on lensing bands: annular regions around the critical curve on the observer’s screen, shaped exclusively by the spacetime geometry. We have specifically focused on an equatorial definition of the first lensing band, defined as the area in the image space in which incident light rays that have intersected the equatorial plane at least twice.
The framework relies only on three inputs: (i) a precise (geometric) definition of the lensing bands; (ii) a detection of a persistent VLBI feature; and (iii) the assumption that (i) and (ii) can be identified. Although the first requirement may sound simple at first glance, defining what a lensed image is can be challenging, as there are various ways to do so. The choice of geometric definition and the status of a detection of a persistent VLBI feature are discussed in Section II.1 and Section II.2, respectively.
As an explicit demonstrations of the framework, we have: (i) specified to a lensing-band definition which counts how many times an incident light ray has crossed the equatorial plane; and (ii) assumed a detection of an associated VLBI feature for which we model a thin ring. Given these choices, the precise exclusion statement which we quantitatively explore can be stated as:
@@
If one assumes a definition of the “photon ring” as the lensed image associated with the above lensing-band definition, and assumes that a VLBI measurement has detected said photon ring, then the precise exclusion statement can be rephrased as:
@@
In Section II.3, we have detailed the explicit implementation of the lensing-band framework. Our implementation encompasses (i) numerical ray tracing in stationary, axisymmetric, and asymptotically flat but otherwise arbitrary spacetimes, (ii) numerical bisection of potentially non-star-convex lensing-band boundaries, (iii) numerical minimization of the overlap of the resulting geometric lensing bands with an observational prior, and (iv) Bayesian parameter estimation. In this context, we also develop a novel free-floating bisection algorithm, the details of which are delegated to Appendix A.
@@
Alongside the conceptual description in this paper, we provide links to the associated computational toolbox. The developed algorithm can also be used to generate adaptive pixel grids to compute high-resolution black-hole images by putting more pixels within the lensing bands, as done in, for example, Ref. [61]. We highlight that, while the present application is entirely focused on the $n=1$ lensing band, our framework is entirely general and can straightforwardly be applied to higher-order lensing bands, see [49] for related work on the second photon ring.
@@
Our work is complementary to previous work based on lensing bands. This includes the null tests of GR proposed for the first [45] and second [41, 44] photon rings. Beyond GR, our work complements previous studies which have focused on spherically-symmetric deviations [62] and on deviations of the largest diameter of the second lensing band [49]. In particular, the width and the angular shape of the lensing band have the potential to impose independent constraints. We find that these can be used to lift degeneracies, for instance, between(i) the black-hole mass and its horizon location; (ii) the asymptotic mass and angular momentum; and (iii) PPN parameters and horizon-scale modifications. Nevertheless, the projected constraints obtained from the first lensing band are conservative, i.e., we make no assumption about the astrophysics. This has advantages and disadvantages. On the one hand, such conservative constraints cannot be impacted by incorrect assumptions on the astrophysics. On the other hand, this means that any correct assumption on the astrophysics is expected to improve the constraints.
@@
Turning the projected constraints obtained in this exploratory study into actual constraints will require a concerted effort. As we have highlighted above, all our quantitative statements crucially rest on whether an observed VLBI feature can be identified and confidently associated with a precise definition of lensed emission. We expect that such confidence is best achieved when inference methods based on hybrid imaging [39], the visibility amplitude [45], and polarization [71] will be combined. It is likely that some systematic uncertainty in our assumptions will remain and we thus view the above only as a first step. In a second step, it is imperative to verify any resulting constraints by means of joint inference on parameters in the geometry and the astrophysics. Any such joint-inference study would amount to a test of our assumptions. We emphasize that the latter statement does not rely on a perfect astrophysical modelling within the joint-inference study. It will thus be most valuable to compare to joint-inference studies beyond-GR based on disk models [56, 48], semi-analytic models [121], and GRMHD [57]. Moreover, the second step of joint inference remains crucial since any added insight on the astrophysics of the plasma surrounding the black hole is expected to tighten the obtained constraints.
@@
With observational earth-based VLBI data becoming more precise and space-based VLBI data becoming available in the future, it is imperative to develop generic theoretical tools that go beyond the mere calculation of the critical curve. In this work, we lay the groundwork to quantitatively constrain arbitrary stationary and axisymmetric spacetimes beyond GR with the first lensed image. This contributes, both, to a larger effort to quantify a geometric map between 4-dimensional near-horizon spacetimes and their 2-dimensional optical image on the celestial sphere, and to a larger effort to simultaneously infer geometry and astrophysics.

### 05 Mar 2024 v2

@@ IV.1.3 A seven-parameter family obtained from the leading-order terms in the KRZ expansion
As detailed above, we focus on the KRZ expansion (see Subsec. IV.1.2) of circular spacetimes (see Subsec. IV.1.1). Furthermore, we make sure to maintain oblate spheroidal coordinates at radial asymptotic infinity and match the leading order asymptotics to the mass $M$, the angular momentum $J$, and the PPN parameters $\beta$ and $\gamma$. The respective explicit derivation within the KRZ expansion is detailed in App. B. These choices result in a seven-parameter family of spacetimes for which the metric functions in Eq. 7 can be written in the following simple form
@@
We use the common shorthand $A=J/M$, $\Sigma=r^{2}+A^{2}\cos^{2}\theta$, and $\Delta=r^{2}-2Mr+A^{2}$ which are also often used to denote Kerr spacetime. In addition, we use the shorthand
@@
  • The parameter $a$ accounts for independent deviations of the horizon area [^11] In the context of the KRZ expansion, see App. B, $a$ is identified with the spin parameter of the background Kerr spacetime. away from the horizon area of a Kerr black hole with equatorial horizon radius $r_{0}$ and asymptotic spin $A$.
@@ IV.2 From size constraints to shape constraints
Figure 6: An application of the lensing-band framework to the seven free parameters of the beyond-GR spacetime given in Eq. 8 and the observer’s inclination. Assuming the observation of a VLBI feature, as discussed in Subsec. II.2, these projected constraints, to one standard deviation, result by demanding it to be completely within the first lensing band of the spacetime considered.
Given the spacetime in Eq. 8 and the lensed-emission region in Subsec. II.2, we employ the lensing-band framework to obtain projected constraints on the underlying free parameters of the circular beyond-Kerr geometry given in Eq. 10.
The full results of the Bayesian parameter estimation, including deviations of all seven free parameters $\Xi=(r_{0},M,A,\beta,\gamma,a,a_{01})$ of the underlying spacetime, is presented as corner plot in Fig. 6. As in the previous section, we choose flat priors on all the parameters with the following bounds: $0.01\leq r_{0}\leq 5$, $-5\leq A/M\leq 5$, $-1\leq\beta\leq 3$, $-1\leq\gamma\leq 5$, $-1\leq a\leq 1$, and $-5\leq a_{01}\leq 5$, and check that the resulting metric obeys the theoretical bounds presented in App. D. As we will see, these flat priors are deformed into preferred regions, for most of these parameters, demonstrating the constraining power of a confident observation of a lensed emission region such as the one discussed in Subsec. II.2.
We discuss several aspects of these results in more detail in the following section. To aid the discussion, we present explicit lensing bands of respective metric deformations and refer to the marginalized posteriors and covariances in Fig. 6.
@@ IV.2.2 The effects of the inclination
As discussed in Sec. III, for the Kerr spacetime, the definition of equatorial lensing bands is ill-defined at edge-on inclination (i.e., at $i=90^{\circ}$), see, e.g., Fig. 7 in Ref. [45]. Hence, we are cautious to interpret results at close to edge-on inclination.
With the above caveat, and similar to the Kerr case in Sec. III, our results suggest a preference for low, i.e., close to face-on, inclination. Specifically, we find $i\doteqdot{32^{\circ}}^{+31^{\circ}}_{-21^{\circ}}\;$ which remains very similar to the posterior for the inclination obtained for the Kerr benchmark test in Sec. III. Once more, the lensing band becomes very thin at high, i.e., close to edge-on, inclination. Hence, it becomes increasingly delicate (or at some point impossible) to fit the observed lensed emission region into the theoretical lensing band. We reiterate that such projected constraints formally rely on a confident observation of lensed emission everywhere within the associated image region.
@@ IV.2.3 Lifting (part of) the degeneracy among spherically symmetric PPN corrections $(\beta,\gamma)$
The PPN corrections ($\beta$ and $\gamma$) enter as spherically-symmetric deformations. Both PPN corrections remain consistent with the values preferred by GR ($\beta_{\text{GR}}=1$ and $\gamma_{\text{GR}}=1$). Moreover, they are correlated which suggests that the strongest projected constraint can actually be placed on the combination $(\beta-\gamma)$, see Fig. 6. This combination enters in the considered spacetime metric via $\Delta_{\beta\gamma}$ and, therefore, affects the $g_{tt}$ component of the metric, see Eq. 8. Several other works have quantified constraints directly in terms of constraints on the functional form of $g_{tt}$ [47, 38, 119, 63]. However, away from spherical symmetry or the critical curve, the lensing behaviour is not only controlled by the $g_{tt}$ component of the metric.
@@
In addition to this broadening/thinning effect, deformations of $\beta$ or $\gamma$, also lead to an increase/decrease of the overall size of the lensing band. Such a deformation in the overall size occurs generically for many spherically-symmetric deformation parameters and has previously been discussed in Ref. [38]. In fact, most constraints placed on deviations from GR to date reduce the geometrical information to the apparent size of the photon ring [48, 56] and are thus, as such, highly degenerate [38]. The degeneracy with the mass $M$, can, of course, be lifted by means of an external mass measurement [14]. However, without resolving geometric information beyond the overall apparent shadow diameter, degeneracies among various possible spherically-symmetric deformation parameters are impossible to avoid [120]. Our results demonstrate that the first lensing band provides an independent way to break (at least some of these) these degeneracies. In particular, the lensing-band framework captures more information than just one overall diameter. Even in spherical symmetry, it captures continuous information encoded in the width of the lensing band. As a result, these degeneracies can (at least partially) be lifted. For instance, the independent lower bounds on $\beta$ and $\gamma$ indicates such a breaking of degeneracy.
@@ IV.2.4 Beyond PPN corrections: strong-field modifications
We have chosen to include the parameter $a_{01}$ in order to compare with results in the literature [48, 49, 56]. These studies assume black-hole uniqueness and therefore focus on $a_{01}$ and other higher-order deviation parameters. In an expansion around radial asymptotic infinity, $a_{01}$ contributes to the next-to-leading order spherically-symmetric deviations once the leading-order PPN parameters (i.e., $\beta$ and $\gamma$) have been fixed to their GR values. The latter are well-constrained at Solar System scales, while the former is not. In this sense, if one assumes black-hole uniqueness, then $a_{01}$ captures the leading deviations which have not already been constrained by other observations.
We find that the lensing-band framework places an independent projected constraint on $a_{01}$, both from above and from below, e.g, for the considered case $a_{01}\doteqdot-0.06^{+1.79}_{-1.40}$. Together with the theoretical bounds (see App. D), the lensing-band shape itself seems to be sufficient to break the degeneracy with all other deviation parameters investigated here. On the one hand, a theoretical constraint on the existence of a Killing horizon (cf. App. D) bounds the value of $a_{01}$ from below [59]. On the other hand, larger values of $a_{01}$ lead to a thinner lensing band, see Fig. 8, hence, eventually to an upper bound. Contrary to previous results [48, 49, 56], the obtained projected constraint does not rely on an external mass measurement. The projected constraint on $a_{01}$ has the same order of magnitude than the results inferred from the EHT observation of $*{}^{*}$ [34].
@@ IV.2.5 A shape constraint on the asymptotic angular momentum $J$
The lensing-band formalism also provides access to constraints which rely on the shape and not just on the overall size or width of the lensing band. In the following, we refer to these as “shape constraints.” Corresponding variations in the shape of the critical curve and in the second lensing band have been previously considered in Refs. [25, 42]. In particular, deformations of their approximately elliptical shape have lead to a proposed null test of GR [42, 46], see also Ref. [50] for beyond-GR deformations.
An explicit example of such a shape constraint is provided by deformations in the asymptotic angular momentum $J$, see Fig. 9. At large angular momentum, exceeding the viable values in Kerr spacetime, the lensing band shape is deformed so far that the lensed emission region modelled in Subsec. II.2 can no longer be covered by the lensing band. Thus, the underlying geometry is excluded.
@@ V Discussion
We have introduced a framework for leveraging the observation of a lensed image of astrophysical emission surrounding a supermassive black hole to constrain the underlying spacetime geometry systematically. The approach is based on lensing bands: annular regions around the critical curve on the observer’s screen, shaped exclusively by the spacetime geometry (and the observer’s location). We have specifically focused on an equatorial definition of the first lensing band, defined as the area in the image space in which incident light rays that have intersected the equatorial plane at least twice.
The framework relies only on three inputs: (i) a precise (geometric) definition of the lensing bands; (ii) a detection of a persistent VLBI feature; and (iii) the assumption that (i) and (ii) can be identified. Although the first requirement may sound simple at first glance, defining what a lensed image is can be challenging, as there are various ways to do so. The choice of geometric definition and the status of a detection of a persistent VLBI feature are discussed in Subsecs. II.1 and II.2, respectively.
As an explicit demonstrations of the framework, we have: (i) specified to a lensing-band definition which counts how many times an incident light ray has crossed the equatorial plane; and (ii) assumed a detection of an associated VLBI feature for which we model a thin ring. Given these choices, the precise exclusion statement which we quantitatively explore can be stated as follows:
@@
If one assumes a definition of the “photon ring” as the lensed image associated with the above lensing-band definition, and assumes that a VLBI measurement has detected said photon ring, then the precise exclusion statement can be rephrased as follows:
@@
In Subsec. II.3, we have detailed the explicit implementation of the lensing-band framework. Our implementation encompasses (i) numerical ray tracing in stationary, axisymmetric, and asymptotically flat but otherwise arbitrary spacetimes, (ii) numerical bisection of potentially non-star-convex lensing-band boundaries, (iii) numerical minimization of the overlap of the resulting geometric lensing bands with an observational prior, and (iv) Bayesian parameter estimation, see also Fig. 3 for a visual synthesis. In this context, we also develop a novel free-floating bisection algorithm, the details of which are delegated to App. A.
@@
Alongside the conceptual description in this paper, we provide links to the associated computational toolbox. The developed algorithm can also be used to generate adaptive pixel grids to compute high-resolution black-hole images by putting more pixels within the lensing bands, as done in, for example, Ref. [64]. We highlight that, while the present application is entirely focused on the $n=1$ lensing band, our framework is entirely general and can straightforwardly be applied to higher-order lensing bands, see Ref. [50] for related work on the second photon ring. Constraints from higher-order photon rings will be much tighter because the thickness of the lensing bands (exponentially) decreases with their order. At the same time, higher-order photon rings are more challenging to resolve in observations.
@@
Our work is complementary to previous work based on lensing bands. This includes the null tests of GR proposed for the first Ref. [46] and second Ref. [42, 45] photon rings. Beyond GR, our work complements previous studies which have focused on spherically-symmetric deviations [65] and on deviations of the largest diameter of the second lensing band [50]. In particular, the width and the angular shape of the lensing band have the potential to impose independent constraints. We find that these can be used to lift degeneracies, for instance, between(i) the black-hole mass and its horizon location; (ii) the asymptotic mass and angular momentum; and (iii) PPN parameters and horizon-scale modifications. Nevertheless, the projected constraints obtained from the first lensing band are conservative, i.e., we make no assumption about the astrophysics. This has advantages and disadvantages. On the one hand, such conservative constraints cannot be impacted by incorrect assumptions on the astrophysics. On the other hand, this means that any correct assumption on the astrophysics is expected to improve the constraints. For instance, a fast radial falloff of the emissivity would significantly tighten constraints as the outer edge of the lensing band can be replaced with a new boundary corresponding to a finite radial location at which the intensity is effectively negligible.
@@
Turning the projected constraints obtained in this exploratory study into actual constraints will require a concerted effort. As we have highlighted above, all our quantitative statements crucially rest on whether an observed VLBI feature can be identified and confidently associated with a precise definition of lensed emission. We expect that such confidence is best achieved when inference methods based on hybrid imaging [40], ringing in visibility space [46], and polarization [74] will be combined. It is likely that some systematic uncertainty in our assumptions will remain and we thus view the above only as a first step. In a second step, it is imperative to verify any resulting constraints by means of joint inference on parameters in the geometry and the astrophysics. Any such joint-inference study would amount to a test of our assumptions. We emphasize that the latter statement does not rely on a perfect astrophysical modelling within the joint-inference study. It will thus be most valuable to compare to joint-inference studies beyond-GR based on disk models [59, 49], semi-analytic models [125], and GRMHD [60]. Moreover, the second step of joint inference remains crucial, since any added insight on the astrophysics of the plasma surrounding the black hole is expected to tighten the obtained constraints.
@@
With observational earth-based VLBI data becoming more precise and space-based VLBI data becoming available in the future, it is imperative to develop generic theoretical tools that go beyond the mere calculation of the critical curve. In this work, we propose a concrete framework to quantitatively constrain arbitrary stationary and axisymmetric spacetimes beyond GR with the first indirect image. This contributes, both, to a larger effort to quantify a geometric map between four-dimensional near-horizon spacetimes and their two-dimensional optical image on the celestial sphere, and to a larger effort to simultaneously infer geometry and astrophysics.